OpenStudy (johan14th):

How would you solve for the slope of the line tangent to the curve (y^2)+(xy+1)^3 = 0 at (2,-1) using implicit. the answer should be 3/4 thanks :)

OpenStudy (johan14th):

\[y^2+(xy+1)^3 = 0 \]

OpenStudy (irishboy123):

have you tried doing it? take each term as it comes.

OpenStudy (johan14th):

\[2y\frac{ dy }{ dx } + 3(xy+1)^2 (y+x \frac{ dy }{ dx })=0\]

OpenStudy (johan14th):

\[\frac{ dy }{ dx } = \frac{ -y(3xy^2+3) }{ 2y+x }\]

OpenStudy (johan14th):

this is what i got but when i try to evaluate it at ( 2,1) i cant get the right answer. Is there something i did wrong?

OpenStudy (irishboy123):

the answer is right. i think your tangent is wrong. i've dome it a different way and get \(y' = -\dfrac{3(xy+1)^2y}{2y + 3(xy+1)^2 x}\)

OpenStudy (irishboy123):

yeah, look at your algebra, your first line seems ok

OpenStudy (irishboy123):

\(2y\frac{ dy }{ dx } + 3(xy+1)^2 (y+x \frac{ dy }{ dx })=0\) \(2y\frac{ dy }{ dx } + 3(xy+1)^2 y+ + 3(xy+1)^2 x \frac{ dy }{ dx }=0\)

OpenStudy (johan14th):

okay i think i have discovered my mistake when i moved the equation from the left to the right

OpenStudy (irishboy123):

yes!

OpenStudy (johan14th):

i just saw it when you wrote it. thank you!

OpenStudy (irishboy123):

mp :p